a player kicks a football at an angle of 45 degrees with an initial speed of 20m/s. A second player on the goal line 60m away in the direction of the kick starts running to receive the ball at that instant. Find the speed of the second player with which he should run so as to catch the ball before it hits the ground

Grade:11

2 Answers

Adarsh
6 years ago
Range of football = u2sin2@/g = 40mTime of flight of football = 2.u.sin@/g = 2√2 ssecond player have to cover a distance of 20 meter in 2√2 second .so, Speed of second player = distance / time = 20/2√2 = 5√2 m/s..
Dev Pandey
2 years ago
Time of flight = 2usin(theta)/g
T=2*20*sin45/10
T=2√2sec
Now speed=distance/time
s=60/2√2
s=30/√2
S=15√2m